OpenStudy (anonymous):

A problem states: "There are 7 more girls than boys in a class. There are 23 students in the class in all. How many girls are there in the class?" What are the unknowns in this problem? A. the number of boys and the number of girls in the class B. the number of students in the class C. the number of boys in the class D. the number of girls in the class

OpenStudy (anonymous):

I think it's C but I just wanna be sure.

OpenStudy (amandafae):

Its C

OpenStudy (amity):

Let boys be x,then girls be x+7,x+x+7=23,,,,,2x=16,,,x=8,and hence boys number is 8 ,and girls number isx+7=8+7=15,

OpenStudy (amity):

Let boys be x,then girls be x+7,x+x+7=23,,,,,2x=16,,,x=8,and hence boys number is 8 ,and girls number isx+7=8+7=15,

OpenStudy (amity):

Answer is c and d

OpenStudy (anonymous):

It was A.

OpenStudy (vincent-lyon.fr):

Answer is A, since the question does not mention neither the number of boys nor the number of girls in the class. Both numbers are unknown.